OpenStudy (anonymous):

please check my answer find (f(w)-f(x))/w-x if f(x)=1/x^2 I got (x-w)/(x^2-w2)

OpenStudy (jamesj):

not quite

OpenStudy (anonymous):

How do you solve it?

OpenStudy (jamesj):

\[ \frac{1/w^2 - 1/x^2}{w-x} = \frac{x^2 - w^2}{x^2w^2(w-x)} = -\frac{x+w}{x^2w^2} \]

OpenStudy (jamesj):

k, where do you live?

OpenStudy (anonymous):

ha why do you want to know that?

OpenStudy (jamesj):

Stephenson is my last name as well

OpenStudy (anonymous):

i live in utah

OpenStudy (jamesj):

ok. Maybe we're distant cousins, but very distant. :-)

myininaya (myininaya):

I wish James was my distant cousin. Hes so smart.

OpenStudy (amistre64):

mormon utah?

OpenStudy (amistre64):

im lds ...

OpenStudy (anonymous):

yes mormon utah. that's the only utah there is ha